(a) x+1/2=x+2/3
(b) 2x+3/x=2/3
(c) 3/x=2/1-x
(d) 2-x/3=2x-1/5
(e) 2x-3/4=2/3x
(f) x-1/2=2-x/3
(g) 1-k/3=k+2/2
(h) 2a-3/4=3-a/3
(i) 3/4=t/t+2
(j) 3/4=t/t+2



Answer by stanbon(75887)

(a) x+1/2=x+2/3
Subtract x from both sides to get:
1/2 = 2/3
Therefore no solution
----------------
(b) (2x+3)/x=2/3
Cross multiply:
6x+9 = 3x-6
3x=-15
x=-5
-------------
(c) 3/x=2/(1-x)
Cross multiply
3-3x=2x
3=5x
x=3/5
---------
(d) (2-x)/3=(2x-1)/5
Cross multiply
10-5x=10x-3
15x=13
x=13/15
----------
etc.
The remaining problems are worked similarly.
Cheers,
Stan H.


(e) (2x-3)/4=2/3x
(f) (x-1)/2=(2-x)/3
(g) (1-k)/3=(k+2)/2
(h) (2a-3)/4=(3-a)/3
(i) 3/4=t/(t+2)
(j) 3/4=t/(t+2)